How do i solve [12x - 4/7] = 5 and it's an absolute value problem.
Answer by edjones(8007)

|12x - 4/7| = 5
.
12x - 4/7 = 5
12x=35/7 + 4/7
x=39/7 * 1/12
=39/84 = 13/28
.
12x - 4/7 = -5
12x=4/7 - 35/7
x=-31/7 * 1/12
=-31/84
.
Ed